# A large data semi-global existence and convergence theorem for vacuum Einstein’s equations

### v3

@@ 1.6. Main Idea of the Proof
From a structural standpoint, the Einstein evolution in vacuum (with vanishing cosmological constant $\Lambda=0$), when expressed in a gauge adapted to the underlying geometric dynamics, constitutes a quasi-linear hyperbolic system subject to elliptic constraints. In particular, unlike the Ricci flow [40, 41, 42], which is manifestly parabolic, or even the semi-linear Yang–Mills heat flow [51], the Einstein equations do not admit any global-in-time regular enough coercive functional of geometric origin that could, in principle, act as a Lyapunov quantity. The absence of such a monotone or dissipative functional precludes, in general, any a priori global control of the evolving geometry beyond the small-data regime.
@@
(23) $\frac{\partial g_{ij}}{\partial T}$ $=-2N\Sigma_{ij}-2\Bigl(1-\frac{N}{n}\Bigr)g_{ij}-(\mathscr{L}_{X}g)_{ij},$
$\frac{\partial\Sigma_{ij}}{\partial T}$ $=-(n-1)\Sigma_{ij}-N\Bigl(\mathrm{Ric}_{ij}+\frac{n-1}{n^{2}}g_{ij}\Bigr)+\nabla_{i}\nabla_{j}\Bigl(\frac{N}{n}-1\Bigr)$
(24) $\quad+2N\Sigma_{ik}\Sigma^{k}_{\;j}-\frac{1}{n}\Bigl(\frac{N}{n}-1\Bigr)g_{ij}-(n-2)\Bigl(\frac{N}{n}-1\Bigr)\Sigma_{ij}-(\mathscr{L}_{X}\Sigma)_{ij},$
@@
In this gauge the spatial operator
$g\longmapsto\mathrm{Ric}[g]+\frac{n-1}{n^{2}}g$
is elliptic. Thus, the system (23)–(68) is hyperbolic in its principal part, augmented by elliptic equations controlling the gauge variables $(N,X)$. The evolution of $\Sigma$ is governed by two competing mechanisms:
  • a linear dissipative term $(n-1)\Sigma_{ij}$, reflecting the intrinsic expansion of slices with negative scalar curvature;
  • a quadratic Riccati-type interaction $2N\Sigma_{ik}\Sigma^{k}_{\;j}$, which tends to induce growth.
Additional nonlinear contributions arise through the curvature term $\mathrm{Ric}_{ij}$ and, more subtly, via the lapse deformation factor (N/n)1(N/n)-1, whose control is itself quadratic in $\Sigma$.
To seek control, one may impose a bootstrap assumption asserting smallness of $\Sigma$ and $\mathfrak{T}:=\mathrm{Ric}[g]-\frac{1}{n}\text{Scal}[g]g$ in appropriate Sobolev (scale-invariant) norms and then derive energy estimates for a suitably corrected energy functional (as constructed in [3]) applied to the pair $(\Sigma,\mathfrak{T})$. An alternative construction considers weighted quantities of the form $(e^{\alpha T}\Sigma,e^{\alpha T}\mathfrak{T})$, for a parameter $\alpha\in(0,n-1)$, coupled with initial data prescribed at a large time $T_{0}=a$. Observe, however, that
$e^{\alpha T}\|(\Sigma,\mathfrak{T})\|\leq\mathcal{I}\quad\text{at}\quad T=T_{0}=a\gg 1$
is equivalent to the smallness, in a scale-free norm, of the unweighted fields $(\Sigma,\mathfrak{T})$. Thus, both approaches reduce essentially to small-data analysis.
Consequently, in the case $\Lambda=0$, no global control for large initial data is attainable via currently available techniques, and in fact such control is not expected to hold in general. In the small-data regime of [3], the linear damping dominates the nonlinear terms, yielding global existence and asymptotic stability. In contrast, when the initial data are large, the quadratic Riccati-type growth overtakes the linear decay, rendering the higher-order geometric quantities uncontrollable over long times. This intrinsic conflict between dissipative and self-amplifying contributions—exacerbated by the quadratic dependence of the lapse deformation on the fields—explains the fundamental obstruction to large-data global-in-time control in the hyperbolic Einstein evolution.
When $\Lambda>0$, the structure of the Einstein evolution equations acquires a qualitatively and quantitatively improved feature. Specifically, all nonlinear Riccati-type contributions and curvature couplings of the form
$\mathfrak{T}_{ij}:=\mathrm{Ric}_{ij}+\frac{n-1}{n^{2}}g_{ij}$
appear multiplied by a universal exponentially decaying factor $e^{-T}$. Schematically, the evolution of $\Sigma_{ij}$ may be expressed as
$\frac{\partial\Sigma_{ij}}{\partial T}$ $\approx-(n-1)\Sigma_{ij}-e^{-T}N\Bigl(\mathrm{Ric}_{ij}+\frac{n-1}{n^{2}}g_{ij}\Bigr)+e^{-T}\nabla_{i}\nabla_{j}\Bigl(\frac{N}{n}-1\Bigr)+2e^{-T}N\Sigma_{ik}\Sigma^{k}_{\;j}$
(27) $\quad-\frac{1}{n}e^{-T}\Bigl(\frac{N}{n}-1\Bigr)g_{ij}-(n-2)e^{-T}\Bigl(\frac{N}{n}-1\Bigr)\Sigma_{ij}.$
The exponential damping factor $e^{-T}$ acts as a geometric suppression of nonlinear interactions, which we entitle as the weak null structure in Einstein-$\Lambda$ setting. Recall that for Minkowski space $(\mathbb{R}^{3+1},\eta)$, the null condition of Klainerman [25] imposes algebraic cancellation in quadratic nonlinear terms of the schematic form
$\Box_{\eta}\varphi=\mathcal{N}(\partial\varphi,\partial\varphi),$
ensuring that no product of two “bad” derivatives occurs in characteristic directions. This mechanism prevents the focusing of nonlinear effects along null hypersurfaces and underlies small-data global existence results. More refined versions of such structures, termed weak null conditions, arise for quasi-linear wave equations, notably in the geometric spacetime harmonic gauge formulation of the Einstein vacuum equations [30].
In the present setting, the term $2N\Sigma_{ik}\Sigma^{k}_{\;j}$ plays an analogous role to the nonlinearity $(\partial\varphi)^{2}$ in the wave equation. Without suppression, this Riccati-type interaction would typically lead to finite-time blow-up. However, due to the presence of the exponentially decaying prefactor $e^{-T}$, arising from the positive cosmological constant, the effective strength of the nonlinearity decreases along the evolution, thereby enabling controlled long-time propagation, even for initial data of large amplitude, provided that the initial time slice is chosen at sufficiently large $T_{0}$.
Furthermore, the curvature obstruction tensor $\mathfrak{T}$ also appears multiplied by $e^{-T}$, which permits direct control of the lower-order norms of $\Sigma$ via integration of equation (27). Notably, such control remains valid without requiring smallness assumptions on $\mathfrak{T}$. This should be contrasted with the vacuum case $\Lambda=0$, cf. (24), where no such damping is present and hence uniform-in-time control of $\Sigma$ cannot be established without imposing smallness on $\mathfrak{T}$ at the outset. Let us illustrate the heuristics.
For $\Lambda>0$, the obstruction tensor $\mathfrak{T}$ approximately satisfies
(28) $\frac{\partial\mathfrak{T}}{\partial T}\approx Ne^{-T}\big(\Delta_{g}\Sigma+\mathrm{Riem}\cdot\Sigma\big)+e^{-T}\nabla^{2}\!\left(\frac{N}{n}-1\right)+e^{-T}\nabla\!\left(\frac{N}{n}-1\right)\Sigma$
(29) $+e^{-T}\mathfrak{T}\cdot\Sigma+\left(\frac{N}{n}-1\right)g.$
In the constant-mean-curvature (CMC) transported spatial gauge ($X=0$), the system of equations (27-28) form a hyperbolic system, supplemented by an elliptic equation for the lapse function $N$. For the sake of exposition, we assume that $N$ can be solved smoothly from this elliptic equation (though the analytic subtleties involved are addressed elsewhere in the introduction).
The analytical difficulty in the $\Lambda>0$ case is to achieve uniform global control of the geometry for all $T\in[T_{0},T_{\infty}]$. In the rescaled setting, this amounts to proving that curvature, deformation, and gauge variables remain uniformly bounded for all future times. To measure these quantities precisely, we introduce three scale-invariant norms:
$\mathcal{O}:=\sum_{I\leq 3}\Big[\|\nabla^{I}\Sigma\|_{L^{2}(M)}+\|\nabla^{I-1}\mathfrak{T}\|_{L^{2}(M)}\Big],\quad\mathcal{F}:=\sum_{I\leq 2}\|e^{\gamma T}\nabla^{I}\Sigma\|_{L^{2}(M)},$
$\mathcal{N}^{\infty}:=\|e^{\gamma T}\Sigma\|_{L^{\infty}(M)}+\Big\|e^{2\gamma T}\!\Big(\frac{N}{n}-1\Big)\Big\|_{L^{\infty}(M)},\qquad\gamma\in(0,1].$
Heuristically:
  • $\mathcal{O}$ controls the curvature energy of the spacetime;
  • $\mathcal{F}$ measures a renormalized “entropy” (reverse) associated to the transverse-traceless part of $\Sigma$;
  • $\mathcal{N}^{\infty}$ quantifies the pointwise amplitude of the gauge and deformation variables.
Together, they form a closed hierarchy of control for the full system.
The proof proceeds by a bootstrap argument, standard in nonlinear geometric analysis. We assume that on some interval $T_{0},T]\subset[T_{0},T_{\infty})$, the following bounds hold:
$\mathcal{O}(T)\leq\Gamma,\qquad\mathcal{F}(T)\leq\mathds{Y},\qquad\mathcal{N}^{\infty}(T)\leq\mathds{L},$
where the bootstrap constants satisfy
$(\mathcal{I}^{0})^{2}+\mathcal{I}^{0}+1<\min\{\Gamma,\mathds{Y},\mathds{L}\},\qquad\Gamma+\mathds{Y}+\mathds{L}\leq e^{T_{0}/5}.$
Starting from large but finite initial data at $T_{0}$,
$\mathcal{O}(T_{0})+\mathcal{F}(T_{0})+\mathcal{N}^{\infty}(T_{0})\lesssim\mathcal{I}^{0},$
the goal is to show that for all future $T$,
$\mathcal{O}(T)+\mathcal{F}(T)+\mathcal{N}^{\infty}(T)\lesssim\mathcal{I}^{0}+1,$
thereby precluding any geometric blow-up and yielding global-in-time control.
The bootstrap closure relies on a hierarchical chain of control:
$\mathcal{N}^{\infty}\quad\rightarrow\quad\mathcal{F}\quad\rightarrow\quad\mathcal{O}.$
The logic is as follows:
@@
    From $\mathcal{N}^{\infty}$ to $\mathcal{F}$. The lapse function NN and the deformation tensor $\Sigma$ are controlled point-wise on each slice via elliptic regularity and Sobolev inequalities. $\mathcal{F}$ is controlled by integration the equation for $\Sigma$. See proposition 4.1 and theorem 4.1 in section 4 for a detailed exposition.
@@
    From $\mathcal{F}$ to $\mathcal{O}$. Once the geometric entropy $\mathcal{F}$ is under control, one derives energy inequalities for the curvature quantities, leading to a uniform bound for $\mathcal{O}$. See the proof of the theorem 4.2 in the section 4.3 for a detailed exposition.
The crucial damping mechanism arises in the mixed term
$\sum_{I\leq 2}e^{-T}\int_{M(T)}N\,\nabla^{I}\mathfrak{T}\,\nabla^{I}\Sigma\,\mu_{g},$
which appears in the $\mathcal{F}$-energy identity (lower order). This term satisfies the bound

### v4

@@ 1.6. Novelty and the main difficulty
To make this precise, recall first the perturbative hypotheses in the work of Fajman–Kröncke [13]. In the expanding negative Einstein case, the background metric $\gamma$ is assumed to satisfy
$\text{Ric}(\gamma)=-\frac{(n-1)}{n^{2}}\gamma,$
and the initial data are required to obey a smallness condition of the form
(14) $\|g_{0}-\gamma\|_{H^{s^{\prime}}}+\|k_{0}+\sqrt{2}\,\gamma\|_{H^{s^{\prime}-1}}<\delta,$
[…]
(15) $(g(T_{0}),K^{TT}(T_{0}))\in B_{\delta}(\gamma_{0},0)\subset H^{s}\times H^{s-1},$
centered at an Einstein background $\gamma_{0}$ (more precisely, at the associated integrable deformation space), together with the shadow gauge condition. In both cases, the theory is perturbative in the strongest possible sense: the initial metric is assumed to be close, in a high Sobolev topology, to a negative Einstein metric, and the transverse–traceless part of the second fundamental form is required to be small in the corresponding Sobolev norm.
The hypothesis of Theorem 1.1 is of a completely different character. One assumes only the uniform metric comparability
$C^{-1}\xi_{0}\leq g_{0}\leq C\xi_{0},$
together with the bound
(16) $\sum_{I=0}^{3}\|\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}+\sum_{I=0}^{2}\Bigl(\|\nabla^{I}\mathfrak{T}[g_{0}]\|_{L^{2}(M)}+\|e^{a}\nabla^{I}\Sigma_{0}\|_{L^{2}(M)}\Bigr)\leq\mathcal{I}^{0},$
where $\mathcal{I}^{0}>0$ is arbitrary and the only smallness requirement is the relative condition
$\mathcal{I}^{0}e^{-a/10}<1.$
In particular, neither $\|\mathfrak{T}[g_{0}]\|_{H^{2}}$ nor the unweighted quantity $\|\Sigma_{0}\|_{H^{3}}$ is required to be small. Thus the present theorem allows a regime in which
$\|\mathfrak{T}[g_{0}]\|_{H^{2}}\sim\mathcal{I}^{0},\qquad\|\Sigma_{0}\|_{H^{3}}\sim\mathcal{I}^{0},$
with $\mathcal{I}^{0}$ as large as one pleases, provided that the initial CMC time $a$ is chosen sufficiently large in terms of $\mathcal{I}^{0}$.
This difference is not merely linguistic; it is quantitative. Indeed, if $\gamma$ is an Einstein metric, then $\mathfrak{T}[\gamma]=0$, and by continuity of the nonlinear map $g\mapsto\mathfrak{T}[g]$ from high Sobolev metrics to curvature in a two–derivative lower Sobolev space, any perturbative hypothesis of the form (14) forces $\|\mathfrak{T}[g_{0}]\|_{H^{s^{\prime}-2}}\lesssim\delta$. Hence an explicitly constructed family with $\|\mathfrak{T}[g_{0}]\|_{H^{2}}$ arbitrarily large eventually lies outside every perturbative neighbourhood covered by the small–data theories. The same is true for the momentum variable: the smallness assumptions in (14) or (15) are incompatible with an unweighted $H^{3}$-norm of $\Sigma_{0}$ of size $\mathcal{I}^{0}\gg 1$.
A second distinction is geometric. The perturbative works start from a negative Einstein background and therefore presuppose the existence of such a metric on the spatial manifold. Theorem 1.1, by contrast, is stated on an arbitrary closed 33-manifold of negative Yamabe type, under the geometric bounds (276)–(277), without assuming the existence of any Einstein metric and without placing the data in a small neighbourhood of any center manifold or moduli space. The result therefore applies to explicitly constructed large data on compact topologies that need not arise as small perturbations of a hyperbolic Einstein geometry.
It is also important to distinguish the present large–data regime from mere smallness of the scalar curvature defect. In dimension 33, the Hamiltonian constraint gives
$R(g_{0})+\frac{2}{3}=|\Sigma_{0}|^{2}.$
Since the $H^{2}$-norm of $\Sigma_{0}$ is weighted by $e^{a}$, one obtains
$\|R(g_{0})+\tfrac{2}{3}\|_{H^{2}}\lesssim\|\Sigma_{0}\|_{H^{2}}^{2}\lesssim e^{-2a}(\mathcal{I}^{0})^{2}.$
Thus the scalar curvature defect may be very small even when $\|\mathfrak{T}[g_{0}]\|_{H^{2}}$ and $\|\Sigma_{0}\|_{H^{3}}$ are very large. This does not place $g_{0}$ in a perturbative neighborhood of an Einstein metric, because the scalar curvature controls only the trace part of the Ricci tensor and does not control the full trace-free Ricci tensor in any comparable topology. The data considered here are therefore genuinely large from the point of view of the full Einstein evolution.
In summary, the present theorem is not a rephrasing of a known small-data stability result at late time. It gives global future control for an explicitly constructed family of initial data satisfying only the relative smallness condition $\mathcal{I}^{0}e^{-a/10}<1$, while permitting arbitrarily large $\|\mathfrak{T}[g_{0}]\|_{H^{2}}$ and arbitrarily large unweighted $\|\Sigma_{0}\|_{H^{3}}$. To the best of our knowledge, this is the first rigorous global large–data convergence theorem for the vacuum Einstein equations with $\Lambda>0$ in CMC gauge on compact spatial manifolds of negative Yamabe type that is not perturbative around a fixed Einstein background.
We next record several additional features of the formulation and of the proof that are essential in the large–data regime.
A first point concerns the choice of gauge. Rather than working with the coupled hyperbolic–elliptic Einstein system in constant mean curvature and spatial harmonic gauge, we work in CMC–transported spatial coordinates, so that the shift vector field vanishes identically, and we formulate the evolution in terms of the pair
$(\Sigma,\mathfrak{T}),\qquad\mathfrak{T}_{ij}:=\text{Ric}_{ij}[g]-\frac{1}{3}R[g]\,g_{ij},$
coupled only to the elliptic lapse equation (83). This point of view is closer in spirit to the classical work of Choquet–Bruhat and York, where one seeks to evolve curvature-type variables together with the second fundamental form. In the present gauge, the metric is recovered from the transport equation (28) along the future-directed normal flow, whereas the higher-order analysis is carried by the evolution system satisfied by $\Sigma$ and $\mathfrak{T}$.
We now make precise why the spatial harmonic gauge is well suited to perturbative theory, but not to the genuinely large–data setting considered here. Let $\xi$ be a fixed smooth background metric on $M$, and let
$V^{i}:=g^{jk}\bigl(\Gamma[g]^{i}_{jk}-\Gamma[\xi]^{i}_{jk}\bigr)$
denote the spatial harmonic gauge vector. Preserving the condition $V=0$ under the Einstein evolution leads to an elliptic equation for the shift vector field $X$ of the form
(17) $\mathcal{P}_{g,\xi}X=\mathcal{F}[g,\Sigma,N],$
where, schematically,
(18) $(\mathcal{P}_{g,\xi}X)^{i}=-\Delta_{g}X^{i}-\text{Ric}[g]^{i}{}_{j}X^{j}+2\nabla^{j}X^{k}\bigl(\Gamma[g]^{i}_{jk}-\Gamma[\xi]^{i}_{jk}\bigr),$
and $\mathcal{F}[g,\Sigma,N]$ depends on the lapse, the second fundamental form, and the metric. This is the operator that must be inverted in order to solve for the shift.
The perturbative theory rests on the fact that when $g$ remains close to a fixed negative Einstein metric $\xi$, the operator $\mathcal{P}_{g,\xi}$ is a small perturbation of the background operator
(19) $\mathcal{P}_{\xi,\xi}X=-\Delta_{\xi}X-\text{Ric}[\xi](X).$
If $\xi$ is Einstein with
$\text{Ric}[\xi]=-\frac{(n-1)}{n^{2}}\xi,$
then
(20) $\mathcal{P}_{\xi,\xi}X=-\Delta_{\xi}X+\frac{(n-1)}{n^{2}}X.$
In particular, for every smooth vector field $X$,
(21) $\int_{M}\langle\mathcal{P}_{\xi,\xi}X,X\rangle_{\xi}\,d\mu_{\xi}$ $=\int_{M}\Bigl(|\nabla^{\xi}X|_{\xi}^{2}+\frac{(n-1)}{n^{2}}|X|_{\xi}^{2}\Bigr)\,d\mu_{\xi}$
$\geq c_{0}\|X\|_{H^{1}(\xi)}^{2},$
for some constant $c_{0}>0$ depending only on $(M,\xi)$. Thus $\mathcal{P}_{\xi,\xi}$ is injective, self-adjoint on $L^{2}(TM,\xi)$, and, by standard elliptic Fredholm theory, an isomorphism
$\mathcal{P}_{\xi,\xi}\colon H^{r+1}(TM,\xi)\longrightarrow H^{r-1}(TM,\xi)$
for every $r\geq 1$. Moreover,
(22) $\|X\|_{H^{r+1}(\xi)}\leq C_{r,\xi}\|\mathcal{P}_{\xi,\xi}X\|_{H^{r-1}(\xi)}.$
The small-data argument is then obtained by perturbing away from $\mathcal{P}_{\xi,\xi}$. Fix $s>\frac{n}{2}+1$. If $g$ satisfies
(23) $\|g-\xi\|_{H^{s}(\xi)}\leq\delta,$
with $\delta$ sufficiently small, then Sobolev multiplication and the smooth dependence of $\Gamma[g]$, $g^{-1}$, and $\text{Ric}[g]$ on $g$ imply the operator bound
(24) $\|(\mathcal{P}_{g,\xi}-\mathcal{P}_{\xi,\xi})X\|_{H^{r-1}(\xi)}\leq C_{r,\xi}\,\|g-\xi\|_{H^{s}(\xi)}\,\|X\|_{H^{r+1}(\xi)},\qquad 1\leq r\leq s.$
In particular, taking $r=1$ and pairing against $X$, one obtains
(25) $\biggl|\int_{M}\bigl\langle(\mathcal{P}_{g,\xi}-\mathcal{P}_{\xi,\xi})X,X\bigr\rangle_{\xi}\,d\mu_{\xi}\biggr|\leq C_{\xi}\,\|g-\xi\|_{H^{s}(\xi)}\,\|X\|_{H^{1}(\xi)}^{2}.$
Combining (21) and (25), we find that if
$C_{\xi}\delta\leq\frac{1}{2}c_{0},$
then
(26) $\int_{M}\langle\mathcal{P}_{g,\xi}X,X\rangle_{\xi}\,d\mu_{\xi}\geq\frac{1}{2}c_{0}\|X\|_{H^{1}(\xi)}^{2}.$
Hence $\mathcal{P}_{g,\xi}$ is injective. Since it is elliptic of index zero and depends continuously on $g$, it follows that
$\mathcal{P}_{g,\xi}\colon H^{r+1}(TM,\xi)\to H^{r-1}(TM,\xi)$
is an isomorphism for all $1\leq r\leq s$, with quantitative estimate
(27) $\|X\|_{H^{r+1}(\xi)}\leq C_{r,\xi}\|\mathcal{P}_{g,\xi}X\|_{H^{r-1}(\xi)}.$
This is the analytic mechanism used in perturbative treatments such as Fajman-Kröncke [13] and Mondal [34]: the shift equation is solvable because the evolving metric remains inside a sufficiently small $H^{s}$-neighbourhood of a fixed negative Einstein background.
The large-data setting of Theorem 1.1 lies outside this framework. First, the theorem does not assume that $M$ carries any fixed negative Einstein metric at all; it assumes only that $M$ is of negative Yamabe type. Second, even when such a metric exists, the initial data constructed in the theorem are not required to satisfy any smallness condition of the form (23). Thus one has no a priori estimate forcing
$\|g(T)-\xi\|_{H^{s}(\xi)}\ll 1$
on the bootstrap interval, and therefore no quantitative perturbative reduction of $\mathcal{P}_{g,\xi}$ to the coercive background operator $\mathcal{P}_{\xi,\xi}$. In particular, the first–order coefficient
$2\nabla^{j}X^{k}\bigl(\Gamma[g]^{i}_{jk}-\Gamma[\xi]^{i}_{jk}\bigr)$
and the zeroth–order Ricci term $\text{Ric}[g](X)$ can no longer be treated as controlled perturbative errors relative to the background geometry. For this reason, the spatial harmonic gauge does not furnish a robust large–data gauge-fixing mechanism for the problem studied here.
This is precisely why we work instead in CMC–transported spatial coordinates. In that gauge the shift vanishes identically, and the analysis is reduced to a hyperbolic evolution system for $(\Sigma,\mathfrak{T})$ coupled to a single elliptic equation for the lapse, whose coefficients are handled directly in terms of the evolving geometry.
A second point is the hierarchy of energy norms used to control the second fundamental form. The top-order energy for $\Sigma$ is uniformly bounded but, in general, does not decay. By contrast, suitably renormalized lower-order norms of $\Sigma$ do decay exponentially. The proof exploits this discrepancy systematically: lower-order decay is first established and then fed back into the top-order energy estimates to control the nonlinear error terms that are borderline at the highest derivative level. This hierarchy is one of the basic mechanisms by which the large-data bootstrap is closed.
Closely related to this is the structure of the nonlinear terms. At the op order, the potentially dangerous interactions are accompanied by an additional factor $e^{-T}$, and are therefore time-integrable. This feature is not a classical null structure; rather, it is an integrable damping mechanism generated by the cosmological expansion. The same phenomenon governs the lapse: although the normalized lapse defect $\frac{N}{n}-1$ need not itself exhibit top-order decay, the terms in which its highest derivatives occur are always accompanied by an $e^{-T}$-weight, whereas the undamped appearances of $\frac{N}{n}-1$ arise only at lower differential order. This separation between top-order weighted terms and lower-order unweighted terms is crucial throughout the bootstrap argument; see Section 1.7.
A further issue is that all Sobolev and elliptic norms are defined relative to the evolving metric $g$. Consequently, one must propagate quantitative control of the geometry of the slices simultaneously with the energy estimates. The metric is governed by the transport equation
(28) $\partial_{T}g_{ij}=-\frac{2\varphi}{\tau}N\Sigma_{ij}-2\Bigl(1-\frac{N}{n}\Bigr)g_{ij}.$
Integrating this identity along the CMC foliation yields control of the metric coefficients and uniform equivalence of the metrics $g(T)$, provided $\Sigma$ and $\frac{N}{n}-1$ are already controlled in the relevant Sobolev norms. What it does not yield is a gain of one additional spatial derivative on $g$. For that reason, one cannot simply freeze the elliptic theory relative to a fixed background metric and appeal to standard coefficient-based regularity estimates at the top order.
The lapse equation illustrates this point. In our normalization it takes the form
(29) $\Delta_{g}\Bigl(\frac{N}{n}-1\Bigr)+\Bigl(|\Sigma|^{2}+\frac{1}{3}\Bigr)\Bigl(\frac{N}{n}-1\Bigr)=-|\Sigma|^{2}.$
If one were to formulate all Sobolev norms relative to a fixed metric $\xi_{0}$, then the usual elliptic theory would require coefficient control at a derivative level that is not directly available from the transport equation for $g$. The resolution is to commute (29) with covariant derivatives $\nabla=\nabla[g]$ of the evolving metric itself. In that formulation, the commutators involve curvature terms rather than uncontrolled higher derivatives of the metric coefficients.
This is precisely where the choice of unknowns becomes decisive. Since the spatial dimension is three, the full Riemann tensor is algebraically determined by the Ricci tensor. Writing
$\text{Ric}[g]=\mathfrak{T}+\frac{1}{3}R[g]\,g$
and using the Hamiltonian constraint
(30) $R[g]+\frac{2}{3}=|\Sigma|^{2},$
one expresses every curvature term appearing in the commuted lapse equation in terms of $\mathfrak{T}$, $\Sigma$, and lower-order contractions thereof. Consequently, the curvature contributions in the elliptic commutators are controlled by the same Sobolev quantities that already enter the hyperbolic part of the argument. This avoids derivative loss and yields the required estimates for the normalized lapse in the same regularity scale as the evolution variables.
Another notable feature of the theorem is the generality of the allowed spatial topology. The argument applies on every smooth, closed, connected, oriented 33-manifold $M$ with non-positive Yamabe invariant $\sigma(M)\leq 0$. Combined with the explicit construction of the initial data class described earlier in the introduction, this places the result well beyond perturbative stability theory near a fixed hyperbolic background. In particular, the theorem is not tied to a distinguished Einstein metric, nor to a center-manifold description of the long-time dynamics.
Finally, the analytic framework developed here appears robust enough to extend beyond the vacuum problem. The combination of expansion-driven integrable damping, lower-order decay, and elliptic estimates formulated relative to the evolving metric should be adaptable to Einstein-$\Lambda$ systems with matter, including, for instance, Einstein-$\Lambda$-Maxwell and Einstein-$\Lambda$-Euler models. Whether one can obtain comparably sharp large-data future-global results in those settings remains an interesting open problem.
@@ 1.7. Main idea of the proof
[…]
(31) $\partial_{T}g_{ij}$ $=-2N\Sigma_{ij}-2\Bigl(1-\frac{N}{n}\Bigr)g_{ij}-(\mathscr{L}_{X}g)_{ij},$
$\partial_{T}\Sigma_{ij}$ $=-(n-1)\Sigma_{ij}-N\Bigl(\text{Ric}_{ij}-\frac{1}{3}R[g]g_{ij}\Bigr)+\nabla_{i}\nabla_{j}\Bigl(\frac{N}{n}-1\Bigr)$
(32) $\qquad+2N\Sigma_{ik}\Sigma^{k}{}_{j}-\frac{1}{n}\Bigl(\frac{N}{n}-1\Bigr)g_{ij}-(n-2)\Bigl(\frac{N}{n}-1\Bigr)\Sigma_{ij}-(\mathscr{L}_{X}\Sigma)_{ij},$
supplemented by the constraint equations
@@
In the spatial harmonic gauge (together with the rescaled Hamiltonian constraint), the linearization of the map
$g\longmapsto\text{Ric}[g]-\frac{1}{3}R[g]g$
is elliptic, so that the gauge variables are determined by elliptic equations on each slice, whereas the pair $(g,\Sigma)$ evolves by a quasilinear hyperbolic system coupled to these elliptic constraints.
The obstruction to large-data global control in (31)-(34) is already visible at the level of the $\Sigma$-equation. The term $(n-1)\Sigma$ is linearly damping, but the remaining terms contain two genuinely nonperturbative mechanisms:
@@
    the curvature defect
    $\mathfrak{T}_{ij}:=\text{Ric}_{ij}-\frac{1}{3}R[g]g_{ij},$
    which measures the deviation from the constant negative sectional curvature geometry (homogeneous and isotropic) singled out by the gauge; and
@@
    the Riccati term $2N\Sigma_{i}{}^{k}\Sigma_{kj}$, which is quadratic and of the same differential order as the damping term.
Moreover, the lapse defect $\frac{N}{n}-1$ is itself determined by an elliptic equation whose source is quadratic in the evolving geometric fields. Thus, even at the level of the formal energy identities, the linearly decaying contribution $(n-1)\Sigma$ is coupled to nonlinear terms that are not accompanied by any time–integrable coefficient. For this reason, every presently available forward stability argument in the $\Lambda=0$ theory is fundamentally perturbative: one closes the estimates only after imposing smallness on the curvature defect and on the trace–free second fundamental form in scale–invariant Sobolev norms. In particular, prescribing the data at a very large initial time $T_{0}=a\gg 1$ does not by itself create a new mechanism, since a bound of the form
$e^{\alpha T_{0}}\|(\Sigma,\mathfrak{T})(T_{0})\|\leq\mathcal{I},\qquad\alpha>0,$
is merely a reformulation of smallness of the unweighted fields at the initial slice.
The situation changes decisively when $\Lambda>0$. After passing to the rescaled CMC variables used throughout the paper, every term in the evolution equations that is potentially dangerous from the point of view of long–time growth appears with an additional coefficient $e^{-T}$. More precisely, the commuted equations may be written schematically as
(35) $\partial_{T}\Sigma$ $=-(n-1)\Sigma+e^{-T}\,\mathcal{Q}_{1}\bigl(g,N,\Sigma,\mathfrak{T},\nabla N,\nabla^{2}N\bigr),$